Tom,Dick,and harry weigh 50,85,and 120lbs.respectively and us a 12ft. plank for a teeterboard.Tom sits 5ft,dick sits 4ft,and harry sits 3ft from the fulcrum on the same sidethe fulcrum is at midpoint of the board.how fare would a 200-pound friend sit from the fulcrum to balance them?
Answer by checkley71(8403)

5*50+4*85+3*120=X*200
250+340+360=200X
950=200X
X=950/200
X=4.75 FT FROM THE FULCRUM FOR THE 200 POUND FRIEND TO SIT & BALANCE THE TEETERBOARD.